# Energy Extraction via Magnetic Reconnection in Kerr-Sen-AdS4 Black Hole: Circular Plasma and Plunging Plasma

### v1

@@ Abstract
This paper investigates the power and efficiency of energy extraction via magnetic reconnection in the Kerr-Sen-AdS4 black hole, focusing primarily on the circular orbit region and the plunging region. After introducing the black hole event horizon, boundary of the ergosphere, and photon sphere radius, we first study the energy extraction via magnetic reconnection in circular orbit region. We plot the allowed regions for energy extraction and present the energy extraction power and efficiency, comparing them with the Blandford-Znajek (BZ) mechanism. It was found that energy extraction is still possible even at a spin parameter of 0.5, which is lower than previous results, and the power can be higher than the BZ process in certain regions. In addition, we find the dilatonic scalar charge $b$ and the AdS radius $l$ contribute to lowering the spin threshold required for energy extraction. We also study the energy extraction in the plunging region by analyzing the allowed regions for energy extraction, and the power as well as efficiency. It was found that the energy extraction is possible even at a spin as low as 0.25. Especially, we find that parameter $b$ promotes lowering the energy extraction spin threshold while $l$ suppresses it, which is different from the circular orbit case. Moreover, the power and efficiency of energy extraction in the plunging region are higher than in the circular orbit region, indicating a stronger energy extraction capability in the plunging region.
@@ I Introduction
Magnetic reconnection is crucial in astrophysics as it rapidly releases magnetic energy, generating significant power. It is often used to explain solar flares 1 , coronal mass ejections 2 and numerous other astronomical phenomena 3 ; 4 . Magnetic reconnection has been extensively studied in magnetohydrodynamics (MHD) 5 ; 6 ; 7 . Recent studies 8 ; 9 ; 10 indicate that spacetime curvature significantly influences the properties of magnetic reconnection, and the effect of gravity on the reconnection process requires further exploration. The analysis of polarized emission around the supermassive black hole in the center of M87* published by the Event Horizon Telescope collaboration 11 also confirms the presence of magnetic fields around black holes. This demonstrates the importance of analyzing magnetic reconnection processes within the framework of general relativity. Furthermore, magnetic reconnection serves as a mechanism for extracting energy from rotating black holes. Comisso and Asenjo first demonstrated in the Kerr black hole that the energy extraction power via this mechanism is typically higher than the traditional Blandford-Znajek mechanism 12 . Subsequently, this work was rapidly extended to other rotating black holes 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 24 ; 25 . However, this mechanism requires the black hole spin to be not too low. For a Kerr black hole, with fixed $\xi=\pi/12,\sigma=100$, the minimum spin is 0.85 12 . In Kerr-de Sitter spacetime, under the same conditions, the minimum spin is no less than 0.7 17 . In rotating regular black hole, under the same conditions and $k=0.1$, the minimum spin is about 0.79 18 . In rotating hairy black hole, under the same conditions and $\lambda=0.5,h_{0}=1.5$, the minimum spin is about 0.86 19 . All the results suggest that the required black hole spin for magnetic reconnection differs for different gravitational backgrounds. Exploring whether energy extraction via magnetic reconnection at lower spins in other gravitational spacetimes thus is a worthwhile topic. In this paper, we study magnetic reconnection in the Kerr-Sen-AdS4 black hole. We intend to investigate how the dilatonic scalar charge $b$ and AdS radius $l$ affect the magnetic reconnection process, and the power and efficiency. Especially whether this system is more effective, compared with previous results, to extract energy via magnetic reconnection. Our results show that the energy extraction is possible even at spins around 0.5 in this system, which is more effective than previous results. This reason stems from that both parameters $b$ and $l$ promote lowering the spin threshold for energy extraction.
@@
Notably, the plasmas in the aforementioned studies all move on circular orbits outside the photon sphere radius. It is worth exploring whether the magnetic reconnection process can be studied in other regions, such as the plunging region. Especially, whether the power and efficiency of energy extraction is the same in different regions. Recently, Ref. 26 extended energy extraction to the plunging region. They found that the energy extraction efficiency in the plunging region is higher than in the circular orbit region, and the allowed region for energy extraction is also larger. In the Kerr-Sen-AdS4 black hole, we also study the magnetic reconnection process in the plunging region, analyzing the impact of the dilatonic scalar charge $b$ and AdS radius $l$ on energy extraction. We find that even at a spin of about 0.25, this black hole can still extract energy in the plunging region, which is lower than Kerr spacetime. We also find that, unlike circular orbits, in the plunging region, $b$ promotes lowering the energy extraction spin threshold, while $l$ suppresses it.

### v2

@@ Abstract
In this paper, we investigate the power and efficiency of energy extraction through magnetic reconnection in the Kerr-Sen-AdS4 black hole, with a primary focus on both the circular orbit and plunging region. We plot the allowed regions for energy extraction and present corresponding power and efficiency values, comparing them with those of the Blandford-Znajek (BZ) mechanism. Our analysis reveals that energy extraction remains feasible even at a spin parameter as low as 0.5, significantly below previously reported thresholds, and the extracted power can exceed that of the BZ process under certain conditions. Furthermore, the dilatonic scalar charge $b$ and the AdS radius $l$ collectively contribute to lowering the spin threshold for energy extraction. The energy extraction process in the plunging region was further examined by analyzing the permissible energy extraction region, as well as the corresponding power output and efficiency. One can find that the energy extraction is possible even at a spin as low as 0.25. Crucially, parameter $b$ actively lowers the energy extraction spin threshold, while parameter $l$ exerts a counteracting effect, impeding such reduction. This trend shows a marked contrast with circular orbit behavior. Additionally, both power and efficiency in the plunging region consistently surpass those in the circular orbit region, indicating superior energy extraction capability in plunging region.
@@ I Introduction
Magnetic reconnection plays a critical role in astrophysics as it enables the rapid release of magnetic energy, thereby generating substantial power. It is frequently employed to account for solar flares 1 , coronal mass ejections 2 , and a wide range of other astronomical phenomena 3 ; 4 . Magnetic reconnection has been extensively studied in magnetohydrodynamics (MHD) 5 ; 6 ; 7 . Recent studies 8 ; 9 ; 10 suggest that spacetime curvature exerts a significant influence on the characteristics of magnetic reconnection, highlighting the need for further investigation into the role of gravitational effects in this process. The analysis of polarized emission surrounding the supermassive black hole at the center of M87*, as reported by the Event Horizon Telescope collaboration (EHT) 11 , provides further confirmation of the existence of magnetic fields in the vicinity of black holes. This highlights the significance of analyzing magnetic reconnection processes within the theoretical framework of general relativity (GR).
On the other hand, magnetic reconnection functions as a key mechanism for extracting energy from rotating black holes. Comisso and Asenjo initially demonstrated that, in the context of a Kerr black hole, the power extracted through this mechanism is generally greater than that achieved via the conventional Blandford-Znajek mechanism 12 . Subsequently, this research was swiftly expanded to encompass other models of rotating black holes 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 24 ; 25 . However, this mechanism requires that the spin of the black hole not be excessively low. For a Kerr black hole with fixed parameters $\xi=\pi/12$ and $\sigma=100$, the minimum spin required is 0.85 12 . In Kerr–de Sitter spacetime, under identical conditions, the minimum spin is found to be no less than 0.7 17 . In a rotating regular black hole, under the same conditions and $k=0.1$, the minimum spin is about 0.79 18 . For a rotating hairy black hole, under the same conditions and $\lambda=0.5,h_{0}=1.5$, the minimum spin is about 0.86 19 . All results suggest that the black hole spin required for magnetic reconnection varies across different gravitational backgrounds. Therefore, exploring the possibility of energy extraction through magnetic reconnection at lower spin values in alternative gravitational spacetimes constitutes a worthwhile research topic. In this paper, we examine magnetic reconnection in the vicinity of the Kerr-Sen-AdS4 black hole. Our objective is to investigate how the dilatonic scalar charge $b$ and the AdS radius $l$ influence the magnetic reconnection process, as well as the associated power and efficiency. Especially, whether this system demonstrates greater effectiveness in extracting energy through magnetic reconnection compared to previous results. We will show that energy extraction is feasible even at a spin value of approximately 0.5 in this system, which represents a more efficient performance compared to previous findings. This improvement can be attributed to the fact that both parameters $b$ and $l$ contribute to lowering the spin threshold required for energy extraction.
@@
It is worth noting that in the aforementioned studies, the plasmas all moved along circular orbits outside the photon sphere radius. Whether magnetic reconnection can be investigated in other regions, such as such as the plunging region, constitutes a highly significant research topic. In particular, whether the power and efficiency of energy extraction remain consistent across different regions. In the work of 26 , the scope of energy extraction has been extended to the plunging region. The results indicate that the efficiency of energy extraction in the plunging region is higher than that in the circular orbit region, and the permissible region for energy extraction is also broader. In the Kerr-Sen-AdS4 black hole spacetime, we investigate the magnetic reconnection process occurring within the plunging region, with a focus on analyzing how the dilatonic scalar charge $b$ and the AdS radius $l$ influence the efficiency of energy extraction. It can be observed that even at a spin parameter of approximately 0.25, this black hole is still capable of extracting energy within the plunging region, a threshold lower than that observed in Kerr spacetime. Furthermore, our findings indicate that, in contrast to circular orbits, the parameter $b$ facilitates a reduction in the spin threshold required for energy extraction within the plunging region, whereas the parameter $l$ has a suppressing effect.